ganeshie8 (ganeshie8):

find all pairs of positive rational numbers \((a,b)\) such that \(\frac{a}{b}+\frac{b}{a}\) is an integer

OpenStudy (freckles):

well an obvious pair would be (1,1)

OpenStudy (empty):

a and b are rational numbers to begin with and not integers?

ganeshie8 (ganeshie8):

yes, \(a\) and \(b\) are rational it seems \((1,1)\) is the only solution...

OpenStudy (empty):

Well what about (2,2) ?

OpenStudy (loser66):

I think set of all a, b such that a =b

OpenStudy (anonymous):

All cases where a=b.

OpenStudy (empty):

;)

OpenStudy (freckles):

\[a=\frac{m}{n} \text{ and } b=\frac{r}{s} \\ \frac{a}{b}+\frac{b}{a}=\frac{(ms)^2+(rn)^2}{mrns}\] I was kinda thinking about doing something with Pythagorean triplets

ganeshie8 (ganeshie8):

Haha right, \(\{(t,t)| t\in \mathbb{Q^{+}}\}\) no other pairs ?

OpenStudy (empty):

Hmmm it would be really nice if we weren't restricted to positive values for a and b.

ganeshie8 (ganeshie8):

this is part of a problem where a, b happen to be positive rationals letting a,b negative simplifies the work is it ?

OpenStudy (empty):

I don't know it just gives more freedom to play around haha. I was thinking of the geometric series and trying to figure out how to make that work. Specifically taking @freckles and putting in something like \[\frac{m^2s^2 + r^2n^2}{mrns} = \frac{x^{2y}-1}{x-1}\] And trying to find solutions here between stuff idk

ganeshie8 (ganeshie8):

Here is the complete problem : https://i.gyazo.com/0c248335ae5de46d24e0ac606c15be22.png

OpenStudy (freckles):

how did you reduce it down to this one

ganeshie8 (ganeshie8):

btw in the actual problem \(a=b\) is not a solution because the left hand side vanishes then..

OpenStudy (freckles):

ah ok

ganeshie8 (ganeshie8):

actually i wanted to show that the only integer value of `a/b + b/a` is 2 when a,b are positive are rationals

OpenStudy (freckles):

oh i see the problem came from @Astrophysics 's post it was the one posted by @mukushla

ganeshie8 (ganeshie8):

yes.. that one..

OpenStudy (freckles):

\[\frac{a}{b}+\frac{b}{a}=2 \\ a^2+b^2=2ab \\ (a-b)^2=0 \\ a=b\]

OpenStudy (freckles):

but a can't be b because that one side vanishes

OpenStudy (freckles):

so that would be mean that one thingy can't be 2

ganeshie8 (ganeshie8):

right, then we conclude that there are no solutions

ganeshie8 (ganeshie8):

it all boils down to showing that \(\frac{a}{b}+\frac{b}{a}\) cannot be an integer when \(a\ne b\)

ganeshie8 (ganeshie8):

\[a=\frac{m}{n} \text{ and } b=\frac{r}{s} \\ \frac{a}{b}+\frac{b}{a}=\frac{(ms)^2+(rn)^2}{mrns} = \dfrac{x^2+y^2}{xy}\] say, \(\dfrac{x^2+y^2}{xy}=t\) \( \implies \color{red}{x}^2-ty\color{red}{x}+y^2=0\) \(\implies x= \dfrac{ty\pm y\sqrt{t^2-4}}{2}\) it follows that \(t^2-4\) must be a perfect square

OpenStudy (empty):

Wait how can \(t^2-4\) be a perfect square when it's \((t-2)(t+2)\)?

ganeshie8 (ganeshie8):

exactly, \(t=2\) is the only positive integer that makes \(t^2-4\) a perfect square (need to prove this tough)

OpenStudy (empty):

\(n^2 \ne x*(x+4)\) sorta seems intuitive for some reason to me because the values are too large and basically consecutive but idk how to reason this out.

OpenStudy (dan815):